OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 6|6 – 4x| = 8x + 4

OpenStudy (anonymous):

6I6-4xI=8x+4 I36-24xI=8x+4 NOW, 36-24x=-(8x+4) AND 36-24x=8x+4

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

ya thank you

OpenStudy (anonymous):

Welcomx

OpenStudy (anonymous):

what dies x =

OpenStudy (anonymous):

??

OpenStudy (anonymous):

x = 1 x = or x = 2 x = x = or x = 1

OpenStudy (anonymous):

thy blank ones are 5 over 2

OpenStudy (anonymous):

36-24x=-(8x+4) AND 36-24x=8x+4 36-24x=-8x-4 -32x=-32 -16x=-40 x=1 x=5/2